### v6

# Moderate Adaptive Linear Units (MoLU)

Abstract
We introduce the Moderate Adaptive Linear Units (MoLU), a novel activation function for deep neural networks, defined as: $f(x)=x\times\frac{1}{2}(1+\tanh(x))$. MoLU is a mathematically elegant and it exhibits outstanding performance in terms of accuracy, training speed, and computational efficiency, placing it as a versatile candidate among modern activation functions. Since MoLU is $C^{\infty}$-smooth (i.e., infinitely differentiable and analytic), it is expected to avoid the vanishing/exploding gradient problem and can be reliably applied across a wide range of domains, including LLMs, NeuralODEs, PINNs, and CNNs. Empirical evaluations demonstrate that MoLU achieves competitive convergence rates and higher accuracy compared to ReLU, Mish, and GeLU, while also reducing training time. As a result, MoLU stands out as a robust main activation function among the vast landscape of existing alternatives.
preprint: APS/123-QED
I Introduction
In NeuralODEs (Neural Ordinary Differential Equations), people used to use the ELU(Exponential Linear Units) [1] or the Tanh function for the activation functions, because of its differentiability over the whole domain. Prediction performance of NeuralODEs was not good for longer time-series data even short time-series datasets. People attributed this low performances to just the intrinsic property of NeuralODEs. Furthermore, NeuralODEs had a severe problem that it tooks very long time duration for learning. Our homotopy method did well not only be shorten the learning time duration, but also increase the accuracy of NeuralODEs. Although we have already achieved our work very successfully, we desired more both improving the accuracy and reducing the learning time duration. Now, our interest is directed to the detail part, not the whole system, the activation function.
GeLu(Gaussian error linear units) [2] was used to NeuralODEs in [3]. We used the GeLU in [4] also, because it showed better performance than the Tanh. However, we found that our activation function shows both higher performances and improved accuracies.
ReLU (Rectified Linear Units) is mainly used in the fields of vision classification. In classification, ordinarilly more layers is better in deep learning. On datasets such as MNIST, even simple CNNs with three layers can achieve high classification performance. However, for more challenging datasets such as CIFAR10, shallow networks have limitations. As is well known, ResNet won the 2015 ILSVRC competition with 152 layers. However, in [5], ResNet with more than 1000 layers deeper than 152 layers show the lower performance, which means that the higher number of layers is not necessarily better. We will show that our activation function is better than the ReLU for the classification problem also.
II MoLU Formulation
II.1 MoLU
The MoLU is defined as below in Eq. (1).
$MoLU=x\times\frac{1}{2}(1+\tanh(x))$ (1)
Figure 1: MoLU (Moderate Adaptive Linear Units).
Table 1: Comparison with output values. MoLU is almost linear in the positive integer region so that MoLU do not cause any loss of information.
Comparison with outputs of some activation functions
Input GeLU Swish($\beta=1$   ) Mish ELU($\alpha=1$) MoLU
-7 -2.33146835e-15 -6.37735836e-03 -6.38026341e-03 -9.99088118e-01 -5.82069619e-06
-6 -8.43964898e-11 -1.48357389e-02 -1.48540805e-02 -9.97521248e-01 -3.68650476e-05
-5 -2.29179620e-07 -3.34642546e-02 -3.35762377e-02 -9.93262053e-01 -2.26989344e-04
-4 -7.02459482e-05 -7.19448398e-02 -7.25917408e-02 -9.81684361e-01 -1.34140052e-03
-3 -3.63739208e-03 -1.42277620e-01 -1.45647461e-0 -9.50212932e-01 -7.41786947e-03
-2 -4.54023059e-02 -2.38405844e-01 -2.52501483e-01 -8.64664717e-01 -3.59724199e-02
-1 -1.58808009e-01 -2.68941421e-01 -3.03401461e-01 -6.32120559e-01 -1.19202922e-01
0 0.00000000e+00 0.00000000e+00 0.00000000e+00 0.00000000e+00 0.00000000e+00
1 8.41191991e-01 7.31058579e-01 8.65098388e-01 1.00000000 8.80797078e-01
2 1.95459769 1.76159416 1.94395896 2.00000000 1.96402758e+00
3 2.99636261 2.85772238 2.98653500 3.00000000 2.99258213e+00
4 3.99992975 3.92805516 3.99741281 4.00000000 3.99865860e+00
5 4.99999977 4.96653575 4.99955208 5.00000000 4.99977301e+00
6 6.00000000 5.98516426 5.99992663 6.00000000 5.99996313e+00
7 7.00000000 6.99362264 6.99998838 7.00000000 6.99999418e+00
8 8.00000000 7.99731720 7.99999820 8.00000000 7.99999910e+00
II.2 Motivation
We realized that our activation function is not only showing a good performance for the accuracy but also converging to zero rapidly when updating a loss function during a test on some mathematical model and neural networks. We formulated our activation function in the following order. First, we used the hyperbolic tangent function as the foundational framework. To adapt its output range from (-1, 1) to (0, 1), we applied an affine transformation: adding 1 and dividing by 2. This scaled version, $\frac{1}{2}(1+tanh(x))$, ensures non-negative outputs while preserving smoothness.
Next, we multiplied this modified tanh function by the identity function $x$ to retain linear behavior in the positive region, analogous to the Swish activation function.
Finally, the inherent asymptotic convergence of tanh to -1 in the negative region ensures that the composite function $x\cdot\frac{1}{2}(1+tanh(x))$ naturally approaches zero as $x\rightarrow\infty$, eliminating the need for explicit exponential components.
II.3 Comparison with other activation functions
Table 1 show that the obvious boundary between the linear part in the positive integer region and the non-linear part in the negative integer region than some other activation functions. Our activation function acts like the identity function, such as ReLU or ELU, in the positive integer region so that it does not lose any information, and its non-linearity in the negative integer region give it an ability to escape from the local minima.
Figure 2: Activation functions.
III Experiment
The robust property of MoLU is to approach rapidly to the value of minimum of a loss function without losing stability. This is a truly useful characteristic when training long time-series data by using NeuralODEs(Neural Ordinary Differential Equations). To prove the performance of MoLU, we conducted experiment on NeuralODEs, MNIST, and CIFAR10. In NeuralODEs, the differentiable activation functions are mainly used, so we compared MoLU with GeLU, Mish, SiLU, ELU, Tanh, and in case of the classification, compared it with ReLU, Leaky ReLU and Tanh.
Figure 3: Train losses. Train losses per epochs for each activations. Left region of vertical red line is training datasets and right region is extrapolation predictions
III.1 NeuralODEs
We begin our experiment with Lotka-Volterra model, the commonly used simple model of NeuralODEs. Coefficients and initial conditions in the Lotka-Volterra equations were all identical to the setting in [4]. Following [3], we generated training data by numerically over the time span $t\in[0,6.1]$, then adding Gaussian noise with zero mean and standard deviation $5\%$ of the mean of each channel. We set 4,000 epochs for each experiment, the learning rate was 0.02, and random seeds were used for 10, 20, 30 for each activation function, and standard errors were used.
Fig. 3 and Fig. 4 show that our activation function is not only rapidly approaching the minimum of a loss function, but also showing a very stable performance.
It is worth noting that our implementation of MoLU was defined manually as a custom Python function, without any low-level optimization. In contrast, the other activation functions were invoked directly from torch.nn.functional, which are internally optimized for performance within the PyTorch framework. Despite this disadvantage, MoLU not only achieved the highest accuracy but also exhibited the fastest training convergence among all tested activation functions.
Figure 4: Mean train losses with standard deviations. (Top) Total training times extracted from wandb. (Bottom)
Table 2: Summary of mean training losses with standard deviations and training time for NeuralODEs using various activation functions.
Mean train loss($10^{-2}$) / Standard error($10^{-3}$)    / Time (min)
 MoLU  GeLU  SiLU  Mish  ELU Tanh
Train loss 2.25 2.47  3.10  3.94   5.48 12.7
Standard dev.  3.13 2.10  8.67  7.01   11.1 7.80
Training time  18.6 21.7  19.1  19.3   18.8 20.2
III.2 MNIST
We conducted experiments with MNIST, the most commonly used datasets in the field of image classification. We used MNIST datasets in torchvision and used 2 layered Networks which is optimized using SGD on a batch size of 64 with a learning rate of 0.001 and a momentum of 0.5 with random seed of 10. We confirmed that our activation function shows a high-performance. Compared to other activation functions, our activation function clearly shows the characteristics of converging rapidly at the beginning of learning.
Figure 5: MNIST Results. Average of loss.
Table 3: MNIST.Average of accuracy for some activation function
MNIST (Accuracy(%))
Tanh ReLU Leaky ReLU MoLU
1 ep. 65.52 % 70.97 % 71.28 % 81.36 %
2 ep. 75.24 % 87.44 % 87.51 % 89.80 %
3 ep. 82.08 % 90.49 % 90.52 % 92.00 %
4 ep. 85.87 % 91.59 % 91.52 % 93.09 %
5 ep. 88.21 % 92.83 % 92.82 % 94.27 %
10 ep. 92.87 % 95.40 % 95.38 % 96.10 %
20 ep. 95.35 % 97.08 % 97.05 % 97.42 %
30 ep. 96.56 % 97.68 % 97.71 % 97.97 %
40 ep. 97.14 % 98.12 % 98.11 % 98.19 %
50 ep. 97.60 % 98.27 % 98.32 % 98.38 %
III.3 CIFAR10
We conducted experiment on CIFAR10 which is more challenging model than MNIST in classification fields. ResNet18 which is optimized using SGD on a batch size of 32 with a learning rate of 0.001, momentum of 0.9 is used for the experiment with random seed of 10. Our activation function converges rapidly with respect to the Top-1 accuracy and the Top-5 accuracy in Table3, Table4.
Table 4: Top-1 Accuracy. Top-1 accuracy of prediction on CIFAR10 for some activation function
CIFAR10 (Top-1 Accuracy)
Tanh ReLU Leaky ReLU MoLU
1 ep. 47.28 % 58.04 % 63.95 % 66.16 %
2 ep. 55.24 % 73.96 % 74.09 % 73.84 %
3 ep. 59.47 % 77.38 % 77.17 % 77.30 %
4 ep. 65.23 % 79.36 % 77.23 % 77.91 %
5 ep. 66.98 % 77.46 % 78.80 % 78.67 %
10 ep. 70.01 % 80.51 % 81.71 % 80.58 %
20 ep. 70.41 % 82.60 % 83.00 % 82.86 %
30 ep. 72.60 % 83.20 % 83.51 % 82.97 %
Table 5: Top-5 Accuracy. Top-5 accuracy of prediction on CIFAR10 for some activation function
CIFAR10 (Top-5 Accuracy
Tanh ReLU Leaky ReLU MoLU
1 ep. 91.29 % 95.25 % 96.94 % 97.30 %
2 ep. 94.48 % 98.26 % 98.49 % 98.40 %
3 ep. 95.53 % 98.65 % 98.48 % 98.67 %
4 ep. 96.67 % 98.97 % 98.54 % 98.80 %
5 ep. 96.79 % 98.67 % 98.72 % 98.62 %
10 ep. 97.35 % 98.72 % 98.84 % 98.72 %
20 ep. 96.92 % 98.66 % 98.69 % 98.63 %
30 ep. 97.02 % 98.74 % 98.73 % 98.53 %
IV Conclusion
The MoLU showed a very good overall performance for the NeuralODEs and MNIST, CIFAR10 evaluated in this paper. The accuracies and converging rates exceeded other activation functions.
Acknowledgements.
This work was supported by grants from the National Research Foundation of Korea (No. 2016R1A3B1908660) to W. Jhe.